OpenStudy (anonymous):

what are the factors of the expression X^2+X-20 @precal

OpenStudy (precal):

ok look at your last number 20

OpenStudy (precal):

list all the factors of 20 (?) times (?)

OpenStudy (precal):

hint 1 times 20 can you think of others?

OpenStudy (anonymous):

1•20 2•10 4•5

OpenStudy (precal):

yes good job, now let's look at the sign in front of 20 what is it?

OpenStudy (anonymous):

negative

OpenStudy (anonymous):

or minus

OpenStudy (precal):

yes and that sign will tell us to subtract the factors, we only want positive numbers so 20-1 =19 we don't want 1-20=-19 can you subtract the other two?

OpenStudy (anonymous):

would that mean they both have to be negative

OpenStudy (precal):

no, both have to be positive

OpenStudy (precal):

20-1=? 10-2=? 5-4=?

OpenStudy (precal):

now keep in mind, if we can factor this, our middle number will appear

OpenStudy (anonymous):

so could you solve it

OpenStudy (precal):

no, I am teaching you how to solve it ever heard of "teach a man to fish, he eats forever, give a man a fish, he eats for one day" You are learning to fish, my friend.......

OpenStudy (precal):

ok so tell me 20-1=19 5-4=what number 10-2 =what number

OpenStudy (anonymous):

10-2

OpenStudy (precal):

10-2=8 5-4=1 what is your middle term?

OpenStudy (anonymous):

Ohhh its 5-4 because there is an invisible 1 in front of the x

OpenStudy (precal):

+1x yes that is correct

OpenStudy (precal):

now we need to figure out our signs

OpenStudy (precal):

a negative sign can only happen if you multiply (+)(-) so we have to figure out the signs of 5 and 4 do we want -5+4 or do we want +5-4 keep in mind we need +1x

OpenStudy (precal):

we are almost done, hang in there

OpenStudy (anonymous):

+5-4

OpenStudy (precal):

yes so the answer is (x+5)(x-4)

OpenStudy (precal):

if you were to multiply this out you would get your original equation

OpenStudy (anonymous):

so then how would i write it in my paper

OpenStudy (precal):

The factors are (x+5)(x-4)

OpenStudy (precal):

just like we did it, I like to show my table of factors

OpenStudy (anonymous):

Because i want to show my work

OpenStudy (precal):

well how does your teacher teach factoring, there are many ways teachers teach factoring

OpenStudy (anonymous):

the usual way a teacher should do it

OpenStudy (precal):

there really is not one way to teach factoring, teachers pick whichever style of doing it. I just showed you my style of factoring

OpenStudy (anonymous):

so how can i write in my paper so that i can show work to it

OpenStudy (precal):